# The 18 May 2024 superbolide over the Iberian Peninsula: USG space sensors and ground-based independent observations

### v2

@@ 3 Results and comparison
The scenario of SPMN180524F originating from a disrupted comet is plausible, as its orbital eccentricity carried the meteoroid well beyond the orbit of Jupiter. The plausible connection between carbonaceous chondrites (CCs) and comets, long suggested, is further supported by recent studies of the reflectance spectra of comet 2P/Encke and ungrouped CCs (Tanbakouei et al., 2020). The role of erosion and subsequent dehydration caused by thermal processing of cometary nuclei was elucidated by Rosetta’s study of 67P/Churyumov-Gerasimenko, providing insights into the evolution of meteoroids resulting from their disruption (Fulle et al., 2020; Koschny et al., 2019; Trigo-Rodríguez et al., 2019). Researches on cometary formation and disintegration products suggest that centimeter-sized pebbles with higher density can be preserved in the interior of comets (Schräpler et al., 2022; Trigo-Rodríguez & Blum, 2022). These pebbles can survive at higher dynamic strengths than the fine-grained materials forming the meteoroid interstitial matrix.
There are examples of superbolides produced by carbonaceous chondrite (CC) projectiles, such as the Tagish Lake meteorite fall, which had an estimated meteoroid density of 1500 kg m-3 but resulted in higher density ungrouped CCs (Brown et al., 2000), as well as the Maribo, Sutter’s Mill, Flensburg, and Winchcombe meteorites (Haack et al., 2012; Jenniskens et al., 2012; Borovička et al., 2021; McMullan et al., 2024). In any case, these fragile meteoroids often generate extremely bright bolides without yielding noticeable meteorites (Borovička et al., 2022). Over more than 25 years of operation, our SPMN network has recorded several cometary superbolides; one of the most memorable events occurred on July 11, 2008, which, due to its anomalous orbit, was tentatively associated with the disruption of comet C/1919 Q2 Metcalf. (Trigo-Rodríguez et al., 2009).
In addition to the physical properties of SPMN180524F, its short perihelion distance suggests a relatively recent formation of the meteoroid. The inferred bulk density is consistent with that expected for a crumbly carbonaceous chondrite. It is well-known that common cometary outgassing cannot release meter-sized meteoroids, indicating that this meteoroid may be a remnant of a relatively recent disruption of its parent comet under the thermal stress imposed by the solar heat near its close perihelion (Jenniskens, 2006; Jewitt, 2008; Trigo-Rodríguez & Blum, 2022).
@@
Efforts have been made to estimate the uncertainties of the CNEOS catalog based on ground-based observations, revealing two groups of measurements: one with sufficient accuracy to allow acceptable heliocentric orbits, and another with significant radiant and velocity deviations Devillepoix et al. (2019) reported discrepancies in the radiants of CNEOS fireballs, ranging from a few degrees to as much as 90. For example, velocity vectors were inaccurately measured for events such as Buzzard Coulee, 2008 TC3, Kalabity, and Crawford Bay. Specific typographical errors, like the missing minus sign in the $z$ velocity component for 2008 TC3, were noted by Peña-Asensio et al. (2022), in addition to comparing two new events independently measured (2019 MO and 2022 EB5). Further independent analyses have included events like Saricicek, Ozerki, Viñales, Flensburg, Novo Mesto, and Adalen, which helped refine the mean radiant and velocity deviations of CNEOS fireballs (Brown & Borovička, 2023; Peña-Asensio et al., 2024).
The superbolide SPMN180524F belongs to the group of events well-measured from space, as the apparent radiant is deviated in 0.98 and the velocity at the energy peak in 1.23 km s-1, resulting in a relatively good agreement on the heliocentric orbit. This difference in velocity leads to a 0.71 au dissent in the semi-major axis. One notable divergence is the 4.43 of disagreement in slope, which yields a -2.81 discrepancy in the orbital inclination. It is worth noting the difference of -288 kg m-3 in density, mainly due to the 3.61 km of disagreement in the height where the radiation peak occurs.
@@ 4 Conclusions
The 18 May 2024 superbolide was a unique event that showcased how a fragile meter-sized meteoroid can produce a spectacular display of color and luminosity, and it exemplified how the Earth’s atmosphere is an excellent shield for this type of impactor. A detailed analysis of the superbolide characteristics suggests that the meteoroid likely originated from the disruption of an eccentric comet, potentially explaining its relatively unstable orbit crossing the orbit of Jupiter. In any case, its short perihelion distance also suggests a recent detachment from its parent comet, as the long-term survival of fragile materials would be challenging under such thermal stress conditions close to the Sun. Given the high velocity and fragmentation observed, a meteorite’s survival seems unlikely. However, the previous Maribo carbonaceous chondrite fall demonstrated that even weak meteoroids can produce small meteorites at relatively high speeds (Borovička et al., 2019).
Our analysis demonstrated relatively good agreement with the data reported by CNEOS both in radiant and velocity, and subsequently in the heliocentric orbital elements. We reconciled the satellite radiometric data with a purely dynamic atmospheric flight model to constrain the meteoroid’s mass and consistently derive the atmospheric deceleration curve.
From an impact risk perspective, the event raises questions about why a meter-sized meteoroid was not detected by current telescopic surveys, especially considering that these small object monitoring programs have successfully identified some asteroids of just a few meters in diameter before their collision with Earth. Our results provide a clear explanation: the meteoroid was simply too small and exhibited a low albedo to be detectable.
@@ Acknowledgements
EP-A acknowledges support from the LUMIO project funded by the Agenzia Spaziale Italiana. JMT-R and PG-T acknowledge financial support from the Spanish project PID2021-128062NB-I00 funded by MCIN/AEI/10.13039/501100011033 . PG-T acknowledges the financial support from the Spanish MCIU through the FPI predoctoral fellowship PRE2022-104624. We also acknowledge the monitoring effort made by the members of the SPMN Network. In particular, Miguel A. Furones, Miguel A. Garcia, and Antonio J. Robles, contributed to the SPMN recordings analyzed in this work. This work is partly based on data from AMS82 cameras of the AllSky7 network, and the authors thank the network operators, especially Juan Carlos Martín, for providing the data.

### v4

@@ 3.1 Atmospheric flight, physical properties, and orbit
The plausible connection between carbonaceous chondrites (CCs) and comets, long suggested, is further supported by recent studies of the reflectance spectra of comet 2P/Encke and ungrouped CCs (Tanbakouei et al., 2020). The role of erosion and subsequent dehydration caused by thermal processing of cometary nuclei was elucidated by Rosetta’s study of 67P/Churyumov-Gerasimenko, providing insights into the evolution of meteoroids resulting from their disruption (Fulle et al., 2020; Koschny et al., 2019; Trigo-Rodríguez et al., 2019). Research on cometary formation and disintegration products suggests that centimeter-sized pebbles with higher density can be preserved in the interior of comets (Blum et al., 2022; Ciarniello et al., 2022; Schräpler et al., 2022; Trigo-Rodríguez & Blum, 2022). However, these materials are too friable to explain the fragments observed trailing the bolide head at aerodynamic pressures of 1 MPa (see Fig. 2). Even carbonaceous bodies have been determined to suffer atmospheric fragmentation at $\sim$0.5 MPa (Brož et al., 2024), with ordinary chondritic surviving greater pressures (Popova et al., 2011). The Almahata Sitta meteorite, a fall associated with the impact of asteroid 2008 TC3, exemplifies a polymict breccia (Bischoff et al., 2010; Jenniskens et al., 2010; Kohout et al., 2010), and recent work suggested that was a polymict C1 chondrite parent body (Bischoff et al., 2022). Similarly, we propose that a good scenario for the SPMN180524F superbolide is that the meter-sized weak carbonaceous meteoroid encountered asteroidal debris, becoming a breccia with higher-strength macroscopic particles in its interior. It is known that common cometary outgassing cannot release meter-sized meteoroids, indicating that the SPMN180524F meteoroid may be a remnant of a disruption of its parent comet under the thermal stress imposed by solar heat near its close perihelion (Jenniskens, 2006; Jewitt, 2008; Trigo-Rodríguez & Blum, 2022). If the SPMN180524F superbolide resulted from an object disruption, additional fragments could reach the Earth.
There are examples of superbolides produced by CC projectiles, such as the Tagish Lake meteorite fall, which had an estimated meteoroid density of 1500 kg m-3 but resulted in higher density ungrouped CCs (Brown et al., 2000), as well as the Maribo, Sutter’s Mill, Flensburg, and Winchcombe meteorites (Haack et al., 2012; Jenniskens et al., 2012; Borovička et al., 2021; McMullan et al., 2024). We also find the polymict breccia hypothesis for the Maribo event plausible, as it had a density of 2000 kg m-3 at entry, with some fragments surviving 3-5 MPa (Borovička et al., 2019). Table 4 lists the potential orbital sources for these events, as computed by NEOMOD. Despite having Tisserand parameters typical of JFCs (or close to it, as in the case of Winchcombe and Tagish Lake), NEOMOD determined that these events most likely did not originate from a JFC source, as opposed to SPMN180524F. CC impactors on JFC-like orbits generally have high entry velocities due to their eccentricity, resulting in their complete disintegration in the atmosphere. In contrast, CCs on other orbital paths tend to enter the atmosphere at slower velocities, increasing their chances of survival. This leads to a bias in our inventory of recovered CCs. The direct video evidence of the meter-sized CC from the SPMN180524F event, which followed a JFC-like orbit, provides a clear example of such high-velocity atmospheric ablation. For instance, the Winchcombe fireball had an entry velocity of only 13.5 km s-1, which facilitated fragments reaching the ground. Table 4 also shows two bright fireballs recorded by the SPMN that were tentatively associated with a JFC (Trigo-Rodríguez et al., 2009; Hughes et al., 2022), as well as the April 13, 2021 bolide observed off the coast of Florida and Grand Bahama Island (Hughes et al., 2022). We used a semi-major axis of 4.2 au because NEOMOD cannot extrapolate the results for the nominal values of SPMN110708 and SPMN130413 (instead of 4.5 and 4.3 au, respectively). However, backward numerical integrations over the past 10,000 years suggest that these superbolides are unlikely to be genuine JFCs (Shober et al., 2024).
Table 4: Most likely sources for superbolides produced by carbonaceous or allegedly JFC projectiles as computed by NEOMOD, along with their Tisserand parameters with respect to Jupiter. The estimated probabilities are shown in parentheses.
Impactor $T_{J}$ 1st source 2nd source 3rd source
SPMN180524F 2.55 JFC (100%)
SPMN110708 2.0 JFC (99.3%) 3:1 (0.4%) 2:1 (0.2%)
SPMN130413 2.3 JFC (97.6%) 5:2 (1.1%) 3:1 (0.6%)
Apr. 13 2021 2.57 JFC (92.0%) 5:2 (5.7%) 3:1 (1.3%)
Tagish Lake 3.52 $\nu_{6}$ (78.6%) 3:1 (20.2%) 5:2 (0.6%)
Maribo 2.95 3:1 (75.4%) $\nu_{6}$ (19.6%) 5:2 (4.3%)
Sutter’s Mill 2.81 5:2 (53.5%) JFC (27.0%) 3:1 (13.1%)
Flensburg 2.89 5:2 (81.6%) 3:1 (10.6%) JFC (4.1%)
Winchcombe 3.12 3:1 (80.8%) $\nu_{6}$ (14.8%) 5:2 (3.7%)
In fact, Shober et al. (2024) performed numerical integration and ran NEOMOD on an extensive fireball dataset to demonstrate that the Jupiter-decoupled orbits of fireballs cannot be attributed to a JFC source, contrary to NEOMOD results or the classical Tisserand parameter classification. The study found that most fireballs with JFC-like orbits are decoupled from Jupiter and exhibit more stable orbits. Specifically, 79-92% of JFC-like meteoroids detected by fireball networks are not subject to frequent Jupiter encounters, and fewer than 5% of all fireballs exhibit dynamics similar to genuine JFCs, suggesting that the likely source is the main belt. Conversely, all genuine JFCs reside on orbits that frequently encounter Jupiter. The decoupling of impacting meteoroids from Jupiter may result from non-gravitational forces, which NEOMOD does not account for. However, non-gravitational forces affecting meteoroids, including those from JFCs, are generally less significant than those affecting comets. The primary non-gravitational force of interest is the Yarkovsky drift, but this effect is slow and insufficient to cause significant decoupling from Jupiter’s orbit compared to the effects of close encounters with Jupiter (Bottke et al., 2000). For the specific case of the SPMN180524F superbolide, we computed the Minimum Orbit Intersection Distance (MOID) relative to Jupiter as 1.51 AU using the method described by Baluev & Mikryukov (2019). This approach yields accurate and rapid results by solving a 16th-order polynomial. This value implies that this impactor with a JFC-like orbit, similar to many identified by Shober et al. (2024), must have been ejected from a body already decoupled from Jupiter’s orbit. The mechanisms responsible for this detachment remain unclear. While secular evolution and non-gravitational forces require extended periods to be effective, frequent close encounters with Jupiter challenge orbital stability. Although the immediate precursor parent body may have evolved from a comet, their orbital dynamics are inconsistent with those of JFCs.
Looking at the bottom-right panel of Figure 6, we can observe a cluster of fireballs, including the SPMN180524F superbolide, separated from the rest and close to several comets. These nearby comets are all Solar and Heliospheric Observatory (SOHO) comets, marked with an orange edge. Also in the vicinity, comets of the Machholz complex can be observed, depicted with a pink edge. All of them with extremely stable orbits over 10,000-year timescales, having MOID with Jupiter larger than 0.5 au (Shober et al., 2024). These events would have been traditionally classified as near-Earth JFCs, despite not having chaotic behaviors, which is indicative of an evolved population. Whether these objects originated on the main belt, the Kuiper belt, or the scattered disk is an open question. Nonetheless, we note that NEOMOD considers SOHO and Machholz comets as JFC with 100% confidence. Based on this and following the classification of Jones et al. (2018), the SPMN180524F superbolide can be termed a sunskirter ($q$ < 0.153 au).
Wiegert et al. (2020) observed a lack of meter-sized bodies with near-Sun perihelia and an excess of millimeter-sized meteoroids. This suggests that near-Sun objects do not fragment into meter-sized pieces but instead disintegrate into millimeter-sized particles. They propose that the disruption of near-Sun asteroids, along with the brightening and destruction processes affecting SOHO comets, occurs through meteoroid erosion, where material is removed by high-speed near-Sun meteoroid impacts. This idea may align with the polymict breccia hypothesis. However, we cannot rule out the possibility that SPMN180524F was a conventional C-type body that was captured into a stable orbit and subsequently experienced volatile sublimation and cracking. This process could have caused it to resemble a weaker object, with only a few interior fragments surviving the thermal processing, which are the pieces observed at the end of its atmospheric flight. In any case, the SPMN180524F event could provide valuable insights into the supercatastrophic disruption mechanisms of sunskirters (Granvik et al., 2016). In future studies, we will conduct a numerical analysis of the dynamic evolution of SPMN180524F to evaluate its stability and gain further insights into its origin.
3.2 Comparison with CNEOS data
@@
Efforts have been made to estimate the uncertainties of the CNEOS database based on ground-based observations, revealing two groups of measurements: one with sufficient accuracy to allow acceptable heliocentric orbits and another one with significant radiant and velocity deviations. Devillepoix et al. (2019) reported discrepancies in the radiants of CNEOS fireballs, ranging from a few degrees to as much as 90. For example, velocity vectors were inaccurately measured for events such as Buzzard Coulee, 2008 TC3, Kalabity, and Crawford Bay. Specific typographical errors, like the missing minus sign in the $z$ velocity component for 2008 TC3, were noted by Peña-Asensio et al. (2022), in addition to comparing two new events independently measured (2019 MO and 2022 EB5). Further independent analyses have included events like Saricicek, Ozerki, Viñales, Flensburg, Novo Mesto, and Adalen, which helped refine the mean radiant and velocity deviations of CNEOS fireballs (Brown & Borovička, 2023; Peña-Asensio et al., 2024).
The last column of Table 2 lists the discrepancy with the values derived from CNEOS data. All the parameters compared were obtained independently, except for mass, which was derived from the radiated energy provided by CNEOS, assuming it is perfectly accurate. The SPMN180524F superbolide belongs to the group of events well-measured from space, as the apparent radiant is deviated in 0.56 and the velocity at the energy peak in 0.1 km s-1, resulting in a good agreement on the heliocentric orbit. One notable divergence is the -1.84 discrepancy in the orbital inclination. Although the CNEOS detection is very close to the atmospheric trajectory, the energy peak coordinate appears offset by $\sim$60 km, despite matching in height. This discrepancy may result from the inherent error in the USG space sensor observations. Nevertheless, there are also challenges in determining the point of maximum brightness from ground-based stations due to highly saturated frames, variations introduced by camera optics, and observation conditions, particularly for distant detections.
@@ 4 Conclusions
The 18 May 2024 superbolide was a unique event, demonstrating how a relatively fragile meter-sized meteoroid can produce a spectacular display of color and luminosity. It also exemplified the Earth’s atmosphere as an excellent shield for this type of impactor. We reconciled the satellite radiometric data with a purely dynamic atmospheric flight model to constrain the meteoroid’s mass and consistently derive the atmospheric velocity profile. The analysis of its characteristics indicates that the most likely source of the carbonaceous meteoroid is the JFC region, aligning with the SOHO comet family, as its sunskirting orbit is decoupled from Jupiter. To explain the presence of fragments surviving pressures of 1 MPa, we hypothesize that the meteoroid could contain collisionally implanted higher-strength macroscopic particles, forming a polymict carbonaceous chondrite object, or it may be a heavily thermally processed C-type asteroid.
Our analysis demonstrated good agreement with the data reported by CNEOS both in radiant and velocity, and subsequently in the heliocentric orbital elements. Additionally, we compared it with the EN fireball catalog and identified similar, but centimeter-sized, events. The presence of meter-sized objects in the vicinity of the Sun provides new constraints on the timescales and characteristics of supercatastophic disruption mechanisms. From an impact risk perspective, this event raises questions about why such an impactor was not detected by current telescopic surveys, which have successfully identified some asteroids of a few meters before their collision with Earth. Our results provide a clear explanation: the meteoroid was too small and had a low albedo, making it hardly detectable.
@@ Acknowledgements
EP-A acknowledges support from the LUMIO project funded by the Agenzia Spaziale Italiana. JMT-R and PG-T acknowledge support from the Spanish project PID2021-128062NB-I00 funded by MCIN/AEI, which sustains the SPMN network. PG-T also acknowledges the FPI predoctoral fellowship PRE2022-104624 from the Spanish MCIU. EP-A and JMT-R conducted this work under the Fundación Seneca project (22069/PI/22), Spain. We also acknowledge the monitoring effort made by the members of the SPMN Network. In particular, Miguel A. Furones, Miguel A. Garcia, and Antonio J. Robles contributed to the recordings analyzed in this work. This work is partly based on data from AMS82 cameras of the AllSky7 network, and we thank the network operators, especially Juan Carlos Martín for providing the data. We thank Denis Vida for his assistance in improving the camera calibrations. We also extend our special thanks to Patrick Shober, William F. Bottke, David Nesvorný, and Rogerio Deienno for their help with the orbit source analysis. We also thank Addi Bischoff, Jürgen Blum, and Giovanni Valsecchi for additional insights.